# Numerical Homogenization by Continuous Super-Resolution

### v1

@@ II-B Super-resolution
Figure 2: The proposed super-resolution model for numerical homogenization. Given a coarse fine element solution represented by $X$ and the corresponding coefficient $A$, we learn a conditional feature extractor using off-the-self neural networks. The extracted feature is then sampled based on the random coordinate map $c$, so that we obtain sampled feature points and grid points. An implicit image function (IIF) takes as input the paired feature and grid points to learn the regression model for super-resolution. To enlarge the receptive field of the feature representation, the extracted feature is reformed by sub-pixel interpolation. The results of coarse- and fine-grained pixel regression are summed together with the sampled LR pixels to obtain the final super-resolved pixels.
@@ III Approach: numerical homogenization by super-resolution
The aim of our approach is to only compute a (not necessarily reasonable) coarse approximation based on the finite element method that is then upscaled to a finer scale by a continuous super-resolution approach without using sophisticated mathematical numerical homogenization tools. From knowledge of a coarse finite element solution (which is cheap to obtain but in general not even a good approximation in a macroscopic sense) and the coefficient $A$, we aim to recover a reasonable solution on a finer scale which contains microscopic information as well. To connect this setup to image-based super-resolution strategies, we represent the coarse finite element solution as a low-resolution (LR) pixel image, the fine-scale solution by an high-resolution (HR) image, and the coefficient as an image on an intermediate scale. We aim to upscale solutions independently of the scale difference: starting from a certain coarse approximation on the scale $h$ with some coefficient that is piecewise constant on a grid consisting of boxes with side length $\varepsilon$, our method should be able to recover an upscaled solution on a chosen scale $\delta$, where typically $h>\varepsilon>\delta$. This approach may be understood as a numerical homogenization strategy in the spirit of the above-mentioned multi-scale approaches, but it acts a posteriori. Its idea of upscaling is more abstract than that of traditional numerical homogenization approaches and does not boil down to modifying basis functions individually. In particular, our approach is not necessarily local and different parts of the image and also features potentially learned from other images are taken into account for a global prediction. This is achieved by a continuous super-resolution model, which is trained on a set of coefficients and corresponding coarse and fine solutions. More details are given in the following subsection.
Our numerical homogenization strategy via continuous super-resolution (NH-CSR) is illustrated in Figure 2. We denote by $X\in\mathbb{R}^{(H+1)\times(H+1)}$ the input low-resolution finite element data, where $H+1$ is the dimensions of its height and width and $H=1/h$. The corresponding coefficient $A\in\mathbb{R}^{E\times E}$ with $E=1/\varepsilon$ is also taken as an additional input for the super-resolution model. Note that $A$ was previously defined as a function. However, since it is assumed to be piecewise constant, it may be represented as a pixel image as well. The overall neural network is a coefficient-guided continuous super-resolution model $f(X,A,\alpha)$ to produce an $\alpha$-times super-resolved finite element solution $\hat{Y}\in\mathbb{R}^{{(\alpha H+1)}\times{(\alpha H+1)}}$, where $\alpha\in\mathbb{R}^{+}$. The complete network consists of three components: 1) global feature extraction, 2) learnable grid sampling, and 3) implicit data point regression, as outlined in detail below.
@@
Intuitively, to achieve arbitrary-scale super-resolution, we first need to have a continuous data representation. The coarse finite element data and the coefficient map are discrete with different data ranges. Hence, we propose a conditional feature extraction using an off-the-shelf image processing backbone to extract the dense feature representation as $F=g(d(X),A)$. We first upsample the coarse finite element data via the bicubic interpolation $d(\cdot)$ to the same size as the coefficient map $A$ and then stack them together as input to learn the dense features $F$. Note that we use the off-the-shelf networks as backbones without using their pre-trained parameters as the finite element data are generated randomly without strong semantic meaning as for natural photos. Using more complex networks can gain some improvements. We will discuss these in Section IV.
III-B Implicit data point regression.
For continuous super-resolution, we are interested in learning a local implicit image function (LIIF) [30],
@@
where $s$ is the bilinear grid sampling operation, $r$ represents the 2D coordinates of the global feature map, and $c$ are the target 2D coordinates. We first apply spatial sampling to the global feature map and its coordinates to localize the neighborhood features, then we use MLP to reconstruct the missing data. Here we use the coordinate residues $s(r,c)-c$ to emphasize the distances between nearest neighbors for prediction.
III-C Gabor based grid sampling
The term in (5) describes a grid-based regression model. Using MLP with ReLUs is biased towards learning low-frequency content [30]. Inspired by the recent development of positional encoding [62] and Fourier feature mapping [35], implicit neural representation shows much better results on high-frequency information prediction. Consider an implicit neural representation function $w_{\theta}:\mathbb{R}^{D_{i}}\rightarrow\mathbb{R}^{D_{o}}$ mapping input coordinates to $D_{o}$ output dimensions with learnable parameters $\theta$. The goal is to map the coordinates to another domain of interest. In general, based on a coordinate vector $y_{m-1}$ in the $(m-1)$-th layer, the consecutive coordinate vector is computed by
@@ III-C Gabor based grid sampling
where $\sigma$ is the nonlinear activation function and $W_{m}$, $b_{m}$ are the weights and biases for the $m$-th layer. In previous works [56, 57], learnable $M$-layer MLPs are utilized for coordinate expansion, including the periodic function $\sigma(x)=\sin(\omega x)$ and the plain ReLU function $\sigma(x)=\max(\omega x,0)$, where $\omega$ denotes the learning parameter for the MLP layer. The problem is that they do not offer explicit multi-resolution representation. In our case, the finite element data naturally preserves a multi-resolution ability, and therefore we propose continuous Gabor wavelets $\phi$ for the coordinate mapping, $\phi(x;\omega_{0},s_{0})=\cos(\omega_{0}x)e^{-(s_{0}x)^{2}}$. It is interesting to point out that when $s_{0}=0$, we have $\phi(x;\omega_{0},s_{0})=\cos(\omega_{0}x)$ which is similar to SIREN in [38]. When $\omega_{0}=0$, we have $\phi(x;\omega_{0},s_{0})=e^{-(s_{0}x)^{2}}$, which is the Gaussian nonlinearity in [63]. We apply two layers of such Gabor nonlinearities to the 2D coordinates, such that we can rewrite (5) as
@@
where $J$ is the output feature vector. Furthermore, we modify the attention mechanism to expand the local feature aggregation. Specifically, we reformalize (7) as a non-local filtering process,
@@
where $Q,K,V$ are the query, key, and value to compute the attention. Further, $D$ is the channel dimension of the local feature embedding. It first calculates the inner product of $Q$ and $K$, adds the relative positional bias obtained by a Gabor based coordinate encoding to the result, and produces an attention matrix normalized by the softmax operation. Intuitively, (8) is the weighted version of (7), which learns the feature correlations and aggregates similar patterns for efficient image reconstruction.
III-D Multi-scale feature sampling.
The key component of the proposed network is to design a grid-based multi-scale feature representation. It takes the extracted feature maps $J\in\mathbb{R}^{k^{2}\times{(H+1)\times(H+1)}}$ (where $k$ is the number of channels) and expands the values to a higher dimension. In the feature domain, we want to upscale the features to the desirable size of $k(H+1)\times k(H+1)$. To achieve that, we first generate a coarse grid $r^{\mathrm{LR}}\in\mathbb{R}^{2\times{(H+1)\times(H+1)}}$ and a fine grid $r^{\mathrm{HR}}\in\mathbb{R}^{2\times{k(H+1)\times k(H+1)}}$. Then we apply a coarse and a fine data regression, that is,
[…]
III-E Losses.
To train our model, we combine an $L1$ loss and our proposed stochastic cosine similarity for optimization. The $L1$ loss is given by $\mathrm{L}_{1}(Y,\hat{Y})=\frac{1}{B}\sum_{i=1}^{B}\|{Y_{i}}-\hat{Y}_{i}\|^{1}_{1}$, where $B$ is the batch size. Note that here $Y_{i}$ and $\hat{Y}_{i}$ are the sampled pixels from predicted and ground truth data, respectively. The problem with this loss is that it takes each data point individually to measure the differences. For a 2D finite element solution, the local region shows similar continuous patterns that need to be collectively optimized. In order to supervise the network to learn data points as a whole set, we get inspiration from S3IM [64] to propose a stochastic cosine similarity score (SCS) $\mathrm{L}_{\mathrm{scs}}$ as follows,
@@ IV-A Implementation Details
Datasets. To generate sufficient data samples for model training and analysis, we randomly generate coefficient maps (independently and identically distributed random values in $\{1,100\}$ on a grid of size $128\times 128$ (which corresponds to $\varepsilon=1/128$)) and use a standard finite element code to obtain the corresponding low-resolution and high-resolution finite element solutions as data (we interpret nodal values as pixels). The resolution of LR data is $33\times 3$ 33 (which corresponds to $h=1/32$), the HR data is $(32\alpha+1)\times(32\alpha+1)$, where $\alpha=2,4,8,16,32,64$. The scaling of alpha by multiples of two and the respective size of the images is classical for finite elements, where the mesh is successively refined by halving the mesh size (i.e., the side length of pixels). For the training, we randomly generate binary coefficient maps as mentioned above. For testing, we have two scenarios:
@@
Parameter setting. The global feature extractor can be any off-the-shelf image encoder. We use residual dense networks (RDN) [65] and Swin [50]. In the following discussion, we use RDN as the default encoder as it balances quality and efficiency. We train NH-CSR using the Adam optimizer with the learning rate $10^{-4}$. The learning rate is halved after 50.000 iterations. The batch size is set to 32, and NH-CSR is trained for 100.000 iterations (about 4 hours) on a PC with one NVIDIA V100 GPU using the PyTorch deep learning platform.
Metrics and evaluation. Mean squared errors (MSE) and mean absolute errors (MAE) measure the average pixel differences between ground truth and estimation. Structural similarities (SSIM) [66] measures the structural closeness of ground truth and estimation. To measure the model complexity, we use FLOPs, memory, and the number of parameters to compare the computational efficiency.
Figure 4: Visual comparison of the coarse finite element solutions. We use the proposed method to apply multi-scale upsampling to the coarse finite solution, both in-distribution (upsampling factor no larger than 16) and out-of-distribution (upsampling factor larger than 16) scenarios. Ours achieve the best visualization than others. As before, the pixel values are multiplied by a factor 8 and shifted to a value range between 0 and 255 to better visualize differences. More precisely, the multiplied values are taken modulo 255.
@@
MetaSR [67] 0.5717 0.7555 0.8902 0.8571 1.0274 0.9004 1.3115 0.9048 1.6404 0.9157 1.9924 0.9201
LIIF [30] 0.5912 0.7493 0.8742 0.8574 0.9523 0.9222 1.2045 0.9329 1.5122 0.9291 1.8445 0.9198
LTE [56] 0.5774 0.7070 0.9044 0.8439 0.9554 0.9199 1.1640 0.9329 1.4442 0.9005 1.7445 0.8948
IDM [68] 0.6152 0.7110 0.9125 0.8442 1.0027 0.9145 1.2204 0.9223 1.5548 0.9012 1.8667 0.8910
LIT [57] 0.5719 0.7524 0.8654 0.8578 0.9246 0.9234 1.1576 0.9335 1.4972 0.9300 1.8056 0.9217
@@ IV-B Comparison of the continuous super-resolution with state-of-the-art methods
To show the efficiency of our proposed method, we compare it with four state-of-the-art continuous super-resolution methods: Meta-SR [67], LIIF [30], LTE [56], LIT [57] and IDM [68]. All approaches are reimplemented using the publicly available codes. We retrained their model on our finite element dataset for fair comparison. In Table I, we report the results of using different methods on both in-distribution and out-of-distribution super-resolution scenarios. We can see that ours achieves the best performance in both MSE and SSIM across different upsampling factors. Note that we use MSE to measure the data fidelity between estimation and the ground truth. SSIM is used to measure the structural closeness which indicates whether super-resolution can preserve the semantic patterns. We can also observe that LIIF and LIT are the most competitive approaches. Meta and LTE have lower performance than others, and IDM also does not perform well as it is built on the pre-trained autoencoder [68]. Compared with others, our model shows consistent significant improvements on MSE by $0.01-0.3$ and SSIM by $0.01-0.03$.
@@
Finally, we are also interested in the frequency analysis on the super-resolution results to see the frequency interpolation. The natural way to describe the high-frequency behavior is through a Fourier transform. To do so, we apply a radial power spectral density (RAPSD) which is based on the assumption that by slicing at any angle of the frequency spectrum, we are able to identify changes in the magnitude spectra. If the model is able to predict the missing high-frequency components, the corresponding energy should match. As shown in the left column of Figure 5, we apply RAPSD to out-of-distribution (32-times and 64-times) SR results in the following steps: First, we convert the SR data to the Fourier domain and slice the frequency spectrum in the clockwise direction to sample the spectral magnitude changing along frequencies ((b) in Figure 5). Then we average the power-vs-frequency results on all testing data and plot it in (c) and (d) in Figure 5. We can see that ours better aligns the trend of energy changes along the frequency bands. On the other hand, LIT introduces many spikes along the high-frequency bands. LIIF, Meta and LTE (green, blue and yellow curves) align first on the low-frequency domain and then introduce errors in the middle- and high-frequency domains.

### v2

@@ II-C Combining mathematical and super-resolution approaches
Figure 2: The proposed super-resolution model for numerical homogenization. (a) shows the overall architecture and (b) shows the detailed operator of LIIF used in coarse and fine LIIF modules. Given a coarse fine element solution represented by $X$ and the corresponding coefficient $A$, we learn a conditional feature extractor using off-the-self neural networks. The extracted feature is then sampled based on the random coordinate map $c$, so that we obtain sampled feature points and grid points. Coarse and fine implicit image functions (IIF) take as input the paired feature and grid points learned by the proposed WIRE operation to form the regression model for super-resolution. The results of coarse- and fine-grained pixel regression are summed together with the sampled LR pixels to obtain the final super-resolved pixels.
@@ III Approach: numerical homogenization by super-resolution
The aim of our approach is to only compute a (not necessarily reasonable) coarse approximation based on the finite element method that is then upscaled to a finer scale by a continuous super-resolution approach without using sophisticated mathematical numerical homogenization tools. From knowledge of a coarse finite element solution (which is cheap to obtain but in general not even a good approximation in a macroscopic sense) and the coefficient $A$, we aim to recover a reasonable solution on a finer scale which contains microscopic information as well. To connect this setup to image-based super-resolution strategies, we represent the coarse finite element solution as a low-resolution (LR) pixel image, the fine-scale solution by a high-resolution (HR) image, and the coefficient as an image on an intermediate scale. We aim to upscale solutions independently of the scale difference: starting from a certain coarse approximation on the scale $h$ with some coefficient that is piecewise constant on a grid consisting of boxes with side length $\varepsilon$, our method should be able to recover an upscaled solution on a chosen scale $\delta$, where typically $h>\varepsilon>\delta$. This approach may be understood as a numerical homogenization strategy in the spirit of the above-mentioned multiscale approaches, but it acts a posteriori. Its idea of upscaling is more abstract than that of traditional numerical homogenization approaches and does not boil down to modifying basis functions individually. In particular, our approach is not necessarily local and different parts of the image and also features potentially learned from other images are taken into account for a global prediction. This is achieved by a continuous super-resolution model, which is trained on a set of coefficients and corresponding coarse and fine solutions. More details are given in the following subsection.
Our numerical homogenization strategy via continuous super-resolution (NH-CSR) is illustrated in Figure 2. We denote by $X\in\mathbb{R}^{(H+1)\times(H+1)}$ the input low-resolution finite element data, where $H+1$ is the dimensions of its height and width and $H=1/h$. The corresponding coefficient $A\in\mathbb{R}^{E\times E}$ with $E=1/\varepsilon$ is also taken as an additional input for the super-resolution model. Note that $A$ was previously defined as a function. However, since it is assumed to be piecewise constant, it may be represented as a pixel image as well. The overall neural network is a coefficient-guided continuous super-resolution model $f(X,A,\alpha)$ to produce an $\alpha$-times super-resolved finite element solution $\hat{Y}\in\mathbb{R}^{{(\alpha H+1)}\times{(\alpha H+1)}}$, where $\alpha\in\mathbb{R}^{+}$. As shown in Figure 2(a), the complete network consists of three main components: 1) global feature extraction (blue block), 2) learnable Gabor wavelet-based grid sampling (The WIRE in light-green blocks), and 3) Multiscale implicit image function (MS-IIF) (olive-green blocks, Figure 2(b)).
@@
Intuitively, to achieve arbitrary-scale super-resolution, we first need to have a continuous data representation. The coarse finite element data and the coefficient map are discrete with different data ranges. Hence, we propose a conditional feature extraction using an off-the-shelf image processing backbone to extract the dense feature representation as $F=g(d(X),A)$ (blue block in Figure 2(a)). We first upsample the coarse finite element data via the bicubic interpolation $d(\cdot)$ to the same size as the coefficient map $A$ and then stack them together as input to learn the dense features $F$ via several convolution operations $g$. Note that we use the off-the-shelf networks as backbones without using their pre-trained parameters as the finite element data are generated randomly without strong semantic meaning as for natural photos. Using more complex networks can gain some improvements. We will discuss these in Section IV.
III-B Gabor wavelet based implicit image function.
For continuous super-resolution, we are interested in learning a local implicit image function (LIIF) [LIIF],
@@
where $s$ is the bilinear grid sampling operation, $r$ represents the 2D coordinates of the global feature map, and $c$ are the target 2D coordinates. We first apply spatial sampling to the global feature map and its coordinates to localize the neighborhood features, then we use MLP to reconstruct the missing data. Here, we use the coordinate residues s(r,c)cs(r,c)-c to emphasize the distances between nearest neighbors for prediction.
The term in (5) describes a grid-based regression model. Using MLP with ReLUs is biased towards learning low-frequency content [LIIF]. Inspired by the recent development of positional encoding [attention] and Fourier feature mapping [NERF], implicit neural representation shows much better results on high-frequency information prediction. Consider an implicit neural representation function $w_{\theta}:\mathbb{R}^{D_{i}}\rightarrow\mathbb{R}^{D_{o}}$ mapping input coordinates to $D_{o}$ output dimensions with learnable parameters $\theta$. The goal is to map the coordinates to another domain of interest. In general, based on a coordinate vector $y_{m-1}$ in the (m1)(m-1)-th layer, the consecutive coordinate vector is computed by
@@
where $\sigma$ is the nonlinear activation function and $W_{m}$, $b_{m}$ are the weights and biases for the $m$-th layer. In previous works [lte, clit], learnable $M$-layer MLPs are utilized for coordinate expansion, including the periodic function $\sigma(x)=\sin(\omega x)$ and the plain ReLU function $\sigma(x)=\max(\omega x,0)$, where $\omega$ denotes the learning parameter for the MLP layer. The problem is that they do not offer an explicit multi-resolution representation. In our case, the finite element data naturally preserves a multi-resolution ability, and therefore we propose continuous Gabor wavelets $\phi$ for the coordinate mapping (light-green blocks in Figure 2(a)), $\phi(x;\omega_{0},s_{0})=\cos(\omega_{0}x)e^{-(s_{0}x)^{2}}$. It is interesting to point out that when $s_{0}=0$, we have $\phi(x;\omega_{0},s_{0})=\cos(\omega_{0}x)$ which is similar to SIREN in [siren]. When $\omega_{0}=0$, we have $\phi(x;\omega_{0},s_{0})=e^{-(s_{0}x)^{2}}$, which is the Gaussian nonlinearity in [gaussian_cor]. We apply two layers of such Gabor wavelet nonlinearities to the 2D coordinates, such that we can rewrite (5) as
@@
where $J$ is the output feature vector. Furthermore, we modify the attention mechanism to expand the local feature aggregation. Specifically, we reformalize (7) as a non-local filtering process (Figure 2(b)),
$\begin{matrix}\!\begin{aligned} &Q=s(W_{\mathrm{Q}}(z)),\ K=s(W_{\mathrm{K}}(z)),\ V=s(W_{\mathrm{V}}(z)),\\ &J=f_{\theta}\left[\mathrm{softmax}\left(\frac{QK^{T}}{\sqrt{D}}+\phi(s(r,c)-c)\right)\times V\right],\end{aligned}\end{matrix}$ (8)
where $Q,K,V$ are the query, key, and value to compute the attention. Further, $D$ is the channel dimension of the local feature embedding. It first calculates the inner product of $Q$ and $K$, adds the relative positional bias obtained by a Gabor wavelet based coordinate encoding to the result, and produces an attention matrix normalized by the softmax operation. Intuitively, (8) is the weighted version of (7), which learns the feature correlations and aggregates similar patterns for efficient image reconstruction.
III-C Multiscale implicit image function (MS-IIF).
A key challenge in continuous super-resolution is how to faithfully interpolate across scales: low-frequency components must be preserved to maintain global consistency, while high-frequency details need to be accurately reconstructed for sharp visualization. A single-scale representation often struggles to balance these two aspects, leading either to overly smooth results (if biased toward low frequencies) or to noisy artifacts (if biased toward high frequencies). To address this, we propose a grid-based multiscale implicit image function (MS-IIF) (coarse and fine LIIF modules in Figure 2(a)), which explicitly disentangles coarse- and fine-scale interpolation. It takes the extracted feature maps $J\in\mathbb{R}^{k^{2}\times{(H+1)\times(H+1)}}$ (where $k$ is the number of channels) and expands the values to a higher dimension. In the feature domain, we want to upscale the features to the desirable size of $k(H+1)\times k(H+1)$. To achieve that, we first generate a coarse grid $r^{\mathrm{LR}}\in\mathbb{R}^{2\times{(H+1)\times(H+1)}}$ and a fine grid $r^{\mathrm{HR}}\in\mathbb{R}^{2\times{k(H+1)\times k(H+1)}}$. Then we apply a coarse and a fine data regression, that is,
$\begin{matrix}\!\begin{aligned} &F^{\mathrm{LR}}=f_{\theta}^{\mathrm{LR}}[s(J,c),\phi(s(r^{\mathrm{LR}},c)-c)],\\ &F^{\mathrm{HR}}=f_{\theta}^{\mathrm{HR}}[s(\mathcal{U}(J),c),\phi(s(r^{\mathrm{HR}},c)-c)],\\ &F=\mathrm{MLP}(F^{\mathrm{LR}}+F^{\mathrm{HR}}).\end{aligned}\end{matrix}$ (9)
@@
. To train our model, we combine an $L1$ loss and our proposed stochastic cosine similarity for optimization. The $L1$ loss is given by $\mathrm{L}_{1}(Y,\hat{Y})=\frac{1}{B}\sum_{i=1}^{B}\|{Y_{i}}-\hat{Y}_{i}\|^{1}_{1}$, where $B$ is the batch size. Note that here $Y_{i}$ and $\hat{Y}_{i}$ are the sampled pixels from predicted and ground truth data, respectively. The problem with this loss is that it takes each data point individually to measure the differences. For a 2D finite element solution, the local region shows similar continuous patterns that need to be collectively optimized. In order to supervise the network to learn data points as a whole set, we get inspiration from S3IM [s3im] to propose a stochastic cosine similarity score (SCS) $\mathrm{L}_{\mathrm{scs}}$ as follows,
@@ IV-A Implementation Details
Datasets. To generate sufficient data samples for model training and analysis, we randomly generate coefficient maps (independently and identically distributed random values in $\{1,100\}$ on a grid of size $128\times 128$ (which corresponds to $\varepsilon=1/128$)) and use a standard finite element code to obtain the corresponding low-resolution and high-resolution finite element solutions as data (we interpret nodal values as pixels). The resolution of LR data is $33\times 33$ (which corresponds to $h=1/32$), the HR data is $(32\alpha+1)\times(32\alpha+1)$, where $\alpha=2,4,8,16,32,64$. The scaling of alpha by multiples of two and the respective size of the images is classical for finite elements, where the mesh is successively refined by halving the mesh size (i.e., the side length of pixels). For the training, we randomly generate binary coefficient maps as mentioned above. For testing, we have two scenarios:
@@
Parameter setting. The global feature extractor can be any off-the-shelf image encoder. We use residual dense networks (RDN) [rdn] and Swin [swinir]. In the following discussion, we use RDN as the default encoder as it balances quality and efficiency. We train NH-CSR using the Adam optimizer with the learning rate $10^{-4}$. The learning rate is halved after 50.000 iterations. The batch size is set to 32, and NH-CSR is trained for 100.000 iterations (about 4 hours) on a PC with one NVIDIA V100 GPU using the PyTorch deep learning platform.
@@
MetaSR [METASR] 0.5717 0.7555 0.8902 0.8571 1.0274 0.9004 1.3115 0.9048 1.6404 0.9157 1.9924 0.9201
LIIF [LIIF] 0.5912 0.7493 0.8742 0.8574 0.9523 0.9222 1.2045 0.9329 1.5122 0.9291 1.8445 0.9198
LTE [lte] 0.5774 0.7070 0.9044 0.8439 0.9554 0.9199 1.1640 0.9329 1.4442 0.9005 1.7445 0.8948
IDM [idm] 0.6152 0.7110 0.9125 0.8442 1.0027 0.9145 1.2204 0.9223 1.5548 0.9012 1.8667 0.8910
LIT [clit] 0.5719 0.7524 0.8654 0.8578 0.9246 0.9234 1.1576 0.9335 1.4972 0.9300 1.8056 0.9217
@@
Metrics and evaluation. Mean squared errors (MSE) and mean absolute errors (MAE) measure the average pixel differences between ground truth and estimation. Structural similarities (SSIM) [ssim] measures the structural closeness of ground truth and estimation. To measure the model complexity, we use FLOPs, memory, and the number of parameters to compare the computational efficiency.
Experimental design and robustness. Our experimental setup is grounded in the theory of numerical homogenization, a general mathematical framework that exploits fine-scale solution features (often referred to as oscillations) within coarse-scale simulations. In this context, testing on finite element data with randomly generated coefficients is a standard and theoretically justified practice, as it effectively captures a broad spectrum of possible variations in the underlying medium. These randomized settings are sufficient to evaluate both the robustness and the generalization capability of the proposed method, thereby addressing the main concerns typically raised about experimental validity. The influence of real-world noisy data is not addressed within this manuscript, as its influence is usually analyzed by running many (similar) numerical experiments, which all could be enhanced using our approach.
Figure 4: Visual comparison of the coarse finite element solutions. We use the proposed method to apply multiscale upsampling to the coarse finite solution, both in-distribution (upsampling factor no larger than 16) and out-of-distribution (upsampling factor larger than 16) scenarios. Ours achieves the best visualization among others. As before, the pixel values are multiplied by a factor 8 and shifted to a value range between 0 and 255 to better visualize differences. More precisely, the multiplied values are taken modulo 255.
@@ IV-B Comparison of the continuous super-resolution with state-of-the-art methods
To show the efficiency of our proposed method, we compare it with four state-of-the-art continuous super-resolution methods: Meta-SR [METASR], LIIF [LIIF], LTE [lte], LIT [clit] and IDM [idm]. All approaches are reimplemented using the publicly available codes. We retrained their model on our finite element dataset for fair comparison. In Table I, we report the results of using different methods on both in-distribution and out-of-distribution super-resolution scenarios. We can see that ours achieves the best performance in both MSE and SSIM across different upsampling factors. Note that we use MSE to measure the data fidelity between estimation and the ground truth. SSIM is used to measure the structural closeness which indicates whether super-resolution can preserve the semantic patterns. We can also observe that LIIF and LIT are the most competitive approaches. Meta and LTE have lower performance than others, and IDM also does not perform well as it is built on the pre-trained autoencoder [idm]. Compared with others, our model shows consistent significant improvements on MSE by 0.010.30.01-0.3 and SSIM by 0.010.030.01-0.03.
@@
Finally, we are also interested in the frequency analysis on the super-resolution results to see the frequency interpolation. The natural way to describe the high-frequency behavior is through a Fourier transform. To do so, we apply a radial power spectral density (RAPSD), which is based on the assumption that by slicing at any angle of the frequency spectrum, we are able to identify changes in the magnitude spectra. If the model is able to predict the missing high-frequency components, the corresponding energy should match. As shown in the left column of Figure 6, we apply RAPSD to out-of-distribution (32-times and 64-times) SR results in the following steps:
First, we convert the SR data to the Fourier domain and slice the frequency spectrum in the clockwise direction to sample the spectral magnitude changing along frequencies ((b) in Figure 6). Then we average the power-vs-frequency results on all testing data and plot it in (c) and (d) in Figure 6. We can see that ours better aligns the trend of energy changes along the frequency bands. On the other hand, LIT introduces many spikes along the high-frequency bands. LIIF, Meta and LTE (green, blue and yellow curves) align first on the low-frequency domain and then introduce errors in the middle- and high-frequency domains.